### v1

@@ Abstract
Text-to-image generation from large generative models like Stable Diffusion, DALLE-2, etc., have become a common base for various tasks due to their superior quality and extensive knowledge bases. As image composition and generation are creative processes where the artists need control over various parts of the images being generated. We find that just adding details about parts in the base text prompt either leads to an entirely different image (e.g., missing/incorrect identity) or the extra part details simply being ignored. To mitigate these issues, we introduce PartCraft, which enables image generation based on fine-grained part-level details specified for objects in the base text prompt. This allows more control for artists and enables novel object compositions by combining distinctive object parts. PartCraft first localizes object parts by denoising the object region from a specific diffusion process. This enables each part token to be localized to the right object region. After obtaining part masks, we run a localized diffusion process in each of the part regions based on fine-grained part descriptions and combine them to produce the final image. All the stages of PartCraft are based on repurposing a pre-trained diffusion model, which enables it to generalize across various domains without training. We demonstrate the effectiveness of part-level control provided by PartCraft qualitatively through visual examples and quantitatively in comparison to the contemporary baselines.  [^1] Project Page: https://rangwani-harsh.github.io/PartCraft
Figure 1: Stable Diffusion vs PartCraft (Ours). PartCraft allows the generation of object images with specified attributes (color, style etc.) of parts for the chosen object in the text prompt. The StableDiffusion baseline generations (left) with part details added to prompt lead to inconsistent outputs compared to proposed PartCraft.
@@ 1 Introduction
Image generation with large generative diffusion models like StableDiffusion [29], DALLE [28], etc., has become prevalent due to their superior quality and extensive knowledge bases. These models are trained on large image-caption datasets and are trained to generate images based on a given text prompt (description). As image composition and creation are creative processes where the artists need control over various parts of the image being generated. However, adding additional details about parts in the text prompt either changes the entire image composition for text-to-image models or ignores the artist’s instructions [8] (warbler in Fig. 1).
@@
There are various works that aim to provide spatial controls to image generation, they allow image generation conditioned on segmentation masks  [4], edge maps [41], bounding boxes [9, 22] etc. Popular methods such as ControlNet [41], GLIGEN [22], etc. require specifying training of these conditional modules, which allows for the controllability of these large generative models. Further, there has been some development of training-free approaches [8, 9] which enable controlled generation of objects by modulating the internal cross-attention activations in the diffusion process. This demonstrates that these pre-trained generative models contain information about the spatial parts of the image, and modulating them effectively can lead to effective image compositions.
@@
However, despite the plethora of works focusing on controllability, the semantic controllability of the image generation is still restricted to specifying details at the object level. The object level details can be restrictive, as often creative designers synthesize object parts (e.g. shirt, trouser etc.) and then compose them [19]. Further, variations in semantic parts are often very distinctive and are used to identify objects [24, 11] uniquely. For example, we usually identify species of birds by their unique beak. Due to this, semantic understanding and recognition of parts have been widely studied as a topic in computer vision [11, 25, 17, 12]. Hence, providing controllability for image synthesis at the object semantic part level can enable a large variety of image compositions [19]. Towards this goal, we introduce PartCraft which provides users with an interface, through which they can select the object in the scene and provide a semantic part-level description with fine-grained details for the object generation .
@@
In PartCraft we first develop a scheme to extract part-level segmentation masks from the StableDiffusion model. For this, we introduce a parallel part diffusion process that generates masks for the object parts. The core idea of the approach is that by forcing the part diffusion model to specifically denoise only the object region in the image, it is possible to get an understanding of the locations of various parts in the object. After composing the object through the part model, the information inside the part diffusion model present in attention maps can be used to generate the masks for various object parts. In the second part, PartCraft utilizes the part masks and description of each part provided by the user. We enable users to provide a highly expressive specification of parts by using a Rich-Text [14] interface, which allows specifying of various styles, colors, etc. for each of the parts.
For generating the final image, taking inspiration from recent studies like Rich-Text Generation [14], Multi-Diffusion [4] etc. we compose the various object parts, by running parallel masked diffusion process for each part while combining them periodically. This combination enables harmonious composition of parts and also makes sure to only locally modify the region corresponding to each part of the object specified by part masks. In the PartCraft pipeline we only make use of the pre-trained StableDiffusion model, and don’t perform any kind of training.
@@
We extensively test proposed PartCraft approach for the zero-shot object part segmentation, which is a very challenging setup in computer vision. We evaluate the part segmentation approach on DeepFashion [23] and CUB200 [38] datasets, where our method significantly outperforms in the baseline StableDiffusion. Further, to evaluate the PartCraft image composition process, we also provide quantitvate and user study results, where our method significantly outperformed the baselines in generating consistent images composed of the described parts Fig. 1. We summarize the core contributions of our paper below:
@@
    We introduce PartCraft, a training-free method that enables the generation of object images by providing fine-grained details for the parts of the object. For example, for a bird we can specify a distinctive description of its beak.
@@
    In PartCraft, we introduce a novel Part-Diffusion process, which localizes and provides masks for object parts, for objects generated by the Base Diffusion model (Fig. 3). To localize object parts, we introduce a segmentation scheme that makes use of the attention maps of the base diffusion and part diffusion process to obtain accurate masks for localized parts .
@@
    In PartCraft after localization, we enable the generation of parts from the pre-trained diffusion model based on the Rich-Text description for the parts provided by the user. PartCraft then composes the image such that all object parts are harmoniously blended by combining localized diffusion path (Fig. 2).
@@ 2 Related Works
Text-2-Image Generative Models. The text-2-image models aim at synthesizing images by following a textual description provided as a prompt. These models have recently become mainstream due to their superior quality and significant knowledge base. This is an outcome of the availability of large-scale image caption datasets [30] and highly parallelized GPU clusters. Almost all kinds of generative models, such as GANs [18], Autoregressive [40], and Diffusion models [28, 29], have shown significant improvements in quality due to training on these large image caption datasets. Among these models, the StableDiffusion [29] models which are based on Diffusion in Latent space are popular due to their open-source nature, which we also utilize for experiments in our work.
Text-2-Image Models for Downstream Tasks. Generative models, in general, have been useful for various downstream tasks, particularly ones based on per-pixel prediction like Segmentation [1, 10], Depth Estimation [6], etc. As layers near the image generation output have features that capture the pixel-pixel relation [6]. With the large-scale text-2-image generative models, these models often perform very competitively [39, 34] to discriminative methods, on tasks like segmentation. However, one commonality among most of these segmentation methods is that they operate at granularity of object or instance level. Text-2-image models here are at an advantage as the usual image captions describe the scene at an object level. In this work, we take a step further exploring the object part level knowledge of these text-2-image diffusion models.
Part Discovery and Segmentation. Part Discovery and Attribution were an integral part of computer vision pipelines classically, as these approaches were robust to view point variations [25, 36, 20]. In deep learning, the unsupervised (self-supervised) approaches for part discovery like SCOPS [17] and Unsup-Parts [11] became popular as they generalize to object parts across categories. In this work, we go one step ahead and operate in zero shot unsupervised part segmentation, where we generate part masks from large Text-2-Image diffusion models.
Controllable Image Generation. Masks along with bounding boxes, edge maps, depth maps etc. have been explored to control the generations of text-2-image diffusion models [22, 41, 5] in addition to text. Further, various other approaches [3, 33, 37, 2, 8] achieve control of image semantics, through modulating diffusion models. Despite this the text-2-image generation control at object part level is unexplored, which we do in our work PartCraft.
3 PartCraft
@@
Figure 2: PartCraft takes input: a base prompt, the selected object $\mathbf{o}$, and a Rich-Text description of the parts. The Part Diffusion process generates the object masks for the specified parts. Then, PartCraft runs a parallel Region Diffusion process to generate attributes of the specified parts, guided by base generation’s intermediate outputs.
@@
In this section, we introduce PartCraft, the method to synthesize objects based on the description of parts of the objects. In PartCraft, we ask the user to specify a base prompt and the token for the object it wants to synthesize parts. Then we provide a rich-text [14] based editor to the user to specify the parts and their description. Our work PartCraft, involves two diffusion steps. a) Part Localization. In the first diffusion step, we get a mask for the object we want to divide into parts. Then, we perform denoising from a U-Net condition on parts to fill the masked region of the object, during which it learns to denoise different object parts. Due to this, the attention maps for various parts start highlighting the correct part region, which we use to extract the part mask. The infilling process is the major contribution of the PartCraft method. b) Part Generation. For generating parts we make combine region specific diffusion process for various parts by iteratively merging them inspired by MuliDiffusion [4], Rich-Text Generation [14] etc. However, till now most of these works have combined the diffusion process for generating objects, for the first time we demonstrate it’s effectiveness in generating object parts. We provide the overview of the PartCraft Generation process in Fig. 2. We below describe the Problem setup followed by the Part Generation and Part Localization setups in detail.
@@ 3.1 Problem Setup
We consider each span of tokens $\mathbf{p}_{i}$ as indicative of describing one part of the object, with attributes $\mathbf{a}_{i}$ describing in appearance. The users are provided with the following ways to describe the parts in terms of their attributes $\mathbf{a}_{i}$ of the object. Our design choices are based on the rich-text generation [14], which allows us to explore various controls for part generation. The
Part Description. The part description (i.e. footnote) is an important attribute that specifies the part of an object; for example, for the part ($\mathbf{p}_{i}$) ’crown’ of a bird, we can specify the attribute ($\mathbf{a}_{i}^{p}$ = ‘a crown of a peacock’). This helps in specifying novel part descriptions, which can lead to artistic novel compositions.
Font Color. The font color helps us specify the exact RGB values for the color attribute $\mathbf{a}_{i}^{c}$, we want for the object part $\mathbf{p}_{i}$. The exact value of RGB allows fine grained control over color of the desired part, whereby just specifying specific colors like ‘brick red’ lead to ignorance by Stable Diffusion [14].
Font Style. The font style is indicative of the specific artistic style $\mathbf{a}_{i}^{s}$ like ‘of Claude Monet’ and ‘of Van Gogh’ when synthesizing images of paintings. This instructs the model to generate a part following a specific style for the given prompt and blend it with other parts of the object. The uniqueness of PartCraft is that it can specify part styles even for objects in paintings and comics, as it’s based on extracting knowledge from large-scale text-2-image models (Sec. 5).
Font Size. The font size controls the relative size of each part in a generation [15]. We use the $\mathbf{a}^{w}_{i}$ to denote the size of each token.
@@ 3.2 Part Localization
We now describe our Part Denoising approach to localize various semantic parts in an object. We first run the base stable diffusion model for the given text prompt and extract the token map $\mathcal{M}_{o}$ for the object $\mathbf{o}$ specified by user, by following the technique of clustering the self-attention maps [27, 14]. We perform this operation at time $t$ less then total time steps for the Diffusion process $T$. After obtaining the binary object mask $\mathcal{M}_{o}$ using attention masks [27, 14], we run two parallel diffusion processes in which one contains input from the text prompt and one contains input from the part prompts $\hat{\mathbf{p}}=[\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{n}]$. One token for each of the part specified in the part text prompt. We denote the $\hat{\mathbf{b}}$ to denote the base text prompt and $\hat{\mathbf{p}}$ to denote the part prompt. For the time $t\leq T_{th}$ (by default we use $T_{th}\approx T/2$) we denoise the U-Net in the object region by denoising it with both the combination of the Part prompt output and Base Prompt Output. Due to this, the Part-Based U-Net gets the information regarding various parts in the object (Fig. 3). We now mathematically define the output noise $\epsilon_{t}$ for the diffusion process with Part-Diffusion below:
$\epsilon_{t}=\alpha*\mathcal{M}_{o}\odot D(x_{t},\hat{\mathbf{p}},t)+(1-\alpha)*\mathcal{M}_{o}\odot D(x_{t},\hat{\mathbf{b}},t)+(1-\mathcal{M}_{o})\odot D(x_{t},\hat{\mathbf{b}},t)$ (1)
here, the $\alpha$ is the hyper-parameter which controls the strength of the part prompt diffusion output, and $D$ is the output of the pre-trained U-Net of a text-2-image diffusion model. On performing some algebra, the above expression can be simplified as the following:
$\epsilon_{t}=\alpha*\mathcal{M}_{o}\odot D(x_{t},\hat{\mathbf{p}},t)+(1-\alpha*\mathcal{M}_{o})\odot D(x_{t},\hat{\mathbf{b}},t)$ (2)
@@
The above denoising process is followed for t steps till 00, to complete one denoising iteration and produce a base image that corresponds to the given prompt $\hat{b}$. Keeping a high $T_{th}$ and low $\alpha$, makes minimal changes in output as if the denoising diffusion was done with original prompt $\hat{b}$ (See Supp. Sec.). With this denoising, we obtain attention masks, from which we extract the localization information of part $\mathbf{p}_{t}$. We now describe the part localization process below.
Token Maps for Parts. We first take the part tokens $\hat{\mathbf{p}}$, which are a concatenation of the part names (e.g. ‘beak crown wings’, etc.), which may not make a meaningful text prompt. Hence, we initialize text embeddings for all these tokens by passing the meaningful text prompt having the following template: “ A photo of $\mathbf{p}_{i}$ of a $\mathbf{o}$ ” where $\mathbf{p}_{i}$ is the object part name and $\mathbf{o}$ is the object name. This serves two purposes: first, it makes the embedding meaningful, and second, it introduces some invariance from the order of part specification in part prompt. These embeddings are then passed as text embeddings to the Part U-Net for denoising. After embedding and running the denoising process, we aggregate across heads and time the self-attention maps (from 32 $\times$ 32 resolution) from both the base and part U-Net diffusion, taking inspiration from works that demonstrate that attention can localize objects [8, 14, 32, 24]. We then perform spectral clustering on these attention maps to form. Based on pixel-pixel similarity, k segmentation maps $\hat{\mathbf{M}}$ (32 $\times$ 32). To attribute these K-segments to the part specified by the user, we aggregate the cross attention of Part U-Net diffusion process. For each token $\mathbf{p}_{j}$, we obtain the cross-attention score as follows:
$\mathbf{\hat{m}}_{j}=\frac{c_{j}}{\sum_{k}c_{k}}$ (3)
where $c_{j}$ is the cross attention score for each token. We proceed by averaging the attention heads to obtain the average cross-attention scores and resize them to 32 $\times$ 32 obtaining $\hat{\mathbf{m}}$. We proceed by removing the start of text (‘<sot>’) token before taking the softmax for cross-attention and then re-normalize it [8]. In other works [27, 14], as the tokens correspond to the image being generated, it’s sure that token maps will be meaningful. However, for parts, it’s not true, as the Part U-Net might not localize some parts. For determining if the part is localized, we propose to look at the max value of cross attention map spatially across the resolution; if we find that the following condition is met, the part is localized:
@@
Here $\delta$ is a hyperparameter, and $K$ is the number of parts. is This condition is robust in finding the localized parts (See Sec. 5 for ablation). For the parts $\mathbf{p}_{i}$, which are localized, we normalize the cross-attention map. We follow a dot-product-based protocol to assign each of the K clusters in the attention map to a part, unlike the average attention protocol in the previous works [27, 14].
$\mathbf{\hat{m}}_{j}=L(j)\frac{\mathbf{\hat{m}}_{j}-\min(\mathbf{\hat{m}}_{j})}{\max(\mathbf{\hat{m}}_{j})-\min(\mathbf{\hat{m}}_{j})}+(1-L(j))\;\hat{\mathbf{m}}_{j}$ (5)
We find that dot product of normalized cross attention scores $\hat{\mathbf{m}_{j}}$ for each token with self-attention masks $\mathbf{\hat{M}}_{j}$ works better, as the attention maps in Part-Unet are noisy. Still, they are often correct for the regions which are localized in one specific area only (Refer Suppl. Sec.). Hence, the dot product protocol favors those maps that are only localized in some areas of the image and don’t have high attention values across all parts of the object. After obtaining the dot product scores for each part we assign $\hat{\mathbf{M}}_{j}$ to $\mathbf{p}_{i}$ with highest scores. The mask for the part $\mathbf{p}_{i}$ is finally given as the following:
@@
Here $\epsilon$ is the hyperparameter which controls the minimum similarity required between the attention mask and the part token. We combine the additional attention masks unassigned to any token, and name them as the background (other) token $\mathbf{M}_{b}$. In the following sections, we evaluate the generated part token masks as zero-shot unsupervised part segmentation (See Fig. 4).

### v2

@@ Abstract
Image composition and generation are processes where the artists need control over various parts of the generated images. However, the current state-of-the-art generation models, like Stable Diffusion, cannot handle fine-grained part-level attributes in the text prompts. Specifically, when additional attribute details are added to the base text prompt, these text-to-image models either generate an image vastly different from the image generated from the base prompt or ignore the attribute details. To mitigate these issues, we introduce PartComposer, a training-free method that enables image generation based on fine-grained part-level attributes specified for objects in the base text prompt. This allows more control for artists and enables novel object compositions by combining distinctive object parts. PartComposer first localizes object parts by denoising the object region from a specific diffusion process. This enables each part token to be localized to the right region. After obtaining part masks, we run a localized diffusion process in each part region based on fine-grained part attributes and combine them to produce the final image. All stages of PartComposer are based on repurposing a pre-trained diffusion model, which enables it to generalize across domains. We demonstrate the effectiveness of part-level control provided by PartComposer through qualitative visual examples and quantitative comparisons with contemporary baselines.
Figure 1: Base Generation (left) Comparison with Methods for Generations with Parts Details (right). PartComposer allows the generation of object images with specified attributes (color, style etc.) of parts for the chosen object in the base text prompt. StableDiffusion and Rich-Text [18] methods with part details either ignore the part instructions or generate inconsistent ojects.
@@ 1 Introduction
Image generation with large generative diffusion models like StableDiffusion [40], DALLE [38], etc., has become prevalent due to their superior quality and extensive world knowledge. These models are trained on large image-caption datasets and are trained to generate images based on a given text prompt (description). As image composition and creation are creative processes where the artists need control over various parts of the image being generated. However, adding additional details for controlling part appearance in the text prompt either changes the generated image entirely or ignores the part instructions [9] (Fig. 1).
@@
Various works aim to provide improved spatial controls to image generation, as they allow image generation conditioned on segmentation masks  [5], edge maps [55], bounding boxes [10, 27] etc. Popular methods such as ControlNet [55], GLIGEN [27], etc. require specifying training of these conditional modules, which allows for the controllability of these large generative models. Further, there has been some development of training-free approaches [9, 10] which enable controlled generation of objects by modulating the internal cross-attention activations in the diffusion process. This demonstrates that these pre-trained generative models contain information about the spatial parts of the image, and modulating them effectively can lead to image compositions.
@@
However, despite several attempts  [6, 55], the semantic controllability of the image generation is still restricted to specifying details at the object level. The object level details can be restrictive, as often creative designers synthesize object parts (e.g. shirt, trouser etc.) and then compose them [23]. Further, variations in semantic parts are often very distinctive and are uniquely used to identify objects [30, 12]. For example, we usually identify species of birds by their unique beak. Due to this, semantic understanding and recognition of parts have been widely studied as a topic in computer vision [12, 31, 21, 13]. Hence, providing controllability for image synthesis at the object semantic part level can enable a large variety of image compositions [23]. Towards this goal, we introduce PartComposer, which provides users with an interface through which they can select the object in the scene and provide a semantic part-level description with fine-grained details for the object generation.
@@
In PartComposer, we develop a scheme to extract part-level localization masks from the Diffusion model. We introduce a parallel part diffusion process that generates masks for the object parts. The core idea of the approach is that by forcing the part diffusion model to specifically denoise only the object region in the image, it is possible to understand the locations of various parts in the object. After denoising just the object through the part model, the information inside the part diffusion model present in attention maps can be used to generate the masks for various object parts. In the following Part Generation stage, PartComposer utilizes the part masks and attributes of each part the user provides. We enable users to provide a highly expressive specification of parts by using a Rich-Text [18] interface, which allows the specifying attributes, such as style, color, etc., for each part. For the final image generation, taking inspiration from recent studies like Rich-Text Generation [18], Multi-Diffusion [5] etc., we compose the various object parts, by running parallel masked diffusion process for each part while combining them periodically into the image. This combination enables harmonious composition of parts, and also masking ensures only local modifications to the regions corresponding to each part of the object specified. In the PartComposer method, we only use the pre-trained StableDiffusion model, making it a generalizable and training-free approach.
@@
We extensively test the proposed PartComposer approach for the zero-shot object part segmentation, which is a challenging setup in computer vision. We evaluate the part segmentation approach on DeepFashion [29] and CUB200 [50] datasets, where our method significantly outperforms the baseline StableDiffusion (SD). Further, to evaluate the PartComposer image composition abilities, we also provide quantitative results along with a user preference study, where our method significantly outperforms the baselines in generating images consistent with the described parts (Fig. 1). We summarize the core contributions of our paper below:
@@
    We introduce PartComposer, a training-free method that enables the generation of object images by using provided fine-grained details for the parts of the object. For example, while generating a bird image we can specify a detailed description of its beak.
@@
    In PartComposer, we introduce a novel Part-Diffusion process, which localizes and provides masks for parts of a base object generated by the Diffusion model (Fig. 3). To localize object parts, we introduce a novel segmentation scheme that uses the attention maps of the base diffusion and part diffusion process to obtain accurate masks for localized parts.
@@
    In PartComposer, after localization, we enable the generation of parts from the pre-trained diffusion model based on the Rich-Text description for the parts provided by the user. PartComposer then composes the image to harmoniously blend all object parts with the background by combining localized diffusion paths (Fig. 2).
@@ 2 Related Works
Text-to-Image Generative Models.
@@
Figure 2: PartComposer takes the input of a base prompt, the selected object $\mathbf{o}$, and a Rich-Text description (  i.e., footnote, colors,   etc.) of the parts. The Part Diffusion process generates the object masks for the specified parts. Then, PartComposer runs a parallel Region Diffusion process to generate attributes of the specified parts, guided by base generation’s intermediate outputs.
The text-to-image models synthesize images by following a textual description provided as a prompt. These models have recently become mainstream due to their superior image generation quality and significant knowledge base. This is an outcome of the availability of large-scale image caption datasets [42] and highly parallelized GPU clusters. Almost all kinds of generative models, such as GANs [22], Autoregressive [53, 41], and Diffusion models [38, 40], have shown significant improvements in quality with training on these large image caption datasets. Among these models, the StableDiffusion (SD) [40] models, based on denoising diffusion in latent space are popular due to their open-source nature, which we also utilize for experiments in our work.
Text-to-Image Models for Downstream Tasks. Generative models, in general, have been useful for various downstream tasks, particularly ones based on per-pixel prediction like Segmentation [1, 11], Depth Estimation [7], etc. As layers near the image generation output have features that capture the pixel-pixel relation [7]. With the large-scale text-to-image generative models, these models often perform very competitively [52, 46] to discriminative methods, on tasks like segmentation. However, one commonality among most of these segmentation methods is that they operate at the granularity of object or instance level. Text-to-image models here are at an advantage as the usual image captions describe the scene at an object level. In this work, we take a step further in exploring the object part-level knowledge of these text-to-image diffusion models.
Part Discovery and Segmentation. Part Discovery and Attribution were an integral part of computer vision pipelines classically, as these approaches were robust to viewpoint variations [31, 48, 24]. In deep learning, the unsupervised (self-supervised) approaches for part discovery like SCOPS [21] and Unsup-Parts [12] became popular as they generalize to object parts across categories. In this work, we go one step ahead and operate in a zero-shot unsupervised part segmentation setting, generating part masks from T2I models.
Controllable Image Generation. Masks, bounding boxes, edge maps, depth maps, etc., have been explored to control the generations of text-to-image diffusion models [27, 55, 6, 14, 36, 35] in addition to text. Further, various other approaches [4, 45, 49, 2, 9, 3, 51, 34] achieve control of image semantics through modulating diffusion models. Despite this, the text-to-image generation control at the object part level is under-explored; a recent work [33] tries to do it in a controlled supervised setup using part-masks. Contrary to that, in our work PartComposer we explore a generalized training-free zero-shot setting.
3 Method
@@
In this section, we introduce PartComposer, our method to synthesize objects based on the description of parts of the objects. In PartComposer, we ask the user to specify a base prompt and the token for the object for which it wants to synthesize parts. Then, we provide the user a rich-text [18] editor (Fig. 2) to specify the parts and their description. PartComposer involves two diffusion steps, a) Part Localization: In the early diffusion stage, we get a mask for the object we want to divide into parts. Then, we perform denoising in later stage from a U-Net condition on parts to fill the masked region of the object, during which it learns to denoise different object parts. Due to this, the attention maps for various parts highlight the correct part region, which we use to extract the part mask. The infilling process is the major contribution of the PartComposer method (Fig, 3). b) Part Generation: For generating parts, we combine region-specific diffusion processes for various parts by iteratively merging them inspired by MuliDiffusion [5], Rich-Text Generation [18] etc. However, till now, most of these works have combined the diffusion process for generating objects; for the first time, we have demonstrated its effectiveness in generating object parts. We provide an overview of the PartComposer pipeline in Fig. 2.
@@ 3.1 Problem Setup
We consider each span of tokens $\mathbf{p}_{i}$ as indicative of describing one part of the object, with attributes $\mathbf{a}_{i}$ describing its overall appearance. Our design choices are based on the rich-text generation [18], and we allow the user to specify the following type of attributes for part generation (Fig. 2):
Part Description (i.e. footnote). It is an important attribute that specifies the part of an object; for example, for the part ($\mathbf{p}_{i}$) ‘crown’ of a bird, we can specify the attribute ($\mathbf{a}_{i}^{p}$ = ‘a crown of a peacock’). This helps specify novel part descriptions, which can lead to artistic novel compositions.
Font Color. It helps us specify the exact RGB values for the color attribute $\mathbf{a}_{i}^{c}$ we want for the object part $\mathbf{p}_{i}$. The exact value of RGB allows fine-grained control over the color of the desired part, whereby just specifying specific colors like ‘brick red’ leads to ignorance by Stable Diffusion [18].
Font Style. The font style is indicative of the specific artistic style $\mathbf{a}_{i}^{s}$ like ‘of Claude Monet’ and ‘of Van Gogh’ when synthesizing images of paintings. This instructs the model to generate a part following a specific style for the given prompt and blend it with other parts of the object (Sec. 5).
Font Size. The font size controls the relative size of each part in a generation [19]. We use the $\mathbf{a}^{w}_{i}$ to denote size.
@@ 3.2 Part Localization
For part localization, we first run the base stable diffusion model for the given text prompt and extract the token map $\mathcal{M}_{o}$ for the object $\mathbf{o}$ specified by the user, by following the technique of clustering the self-attention maps [37, 18]. We obtain the mask using this clustering after denoising between the initial step ($T$) until a threshold time step $T_{th}$. After obtaining the binary object mask $\mathcal{M}_{o}$ using attention masks [37, 18], we run two parallel diffusion processes where one contains input from the base prompt and the other contains input from the part prompts $\hat{\mathbf{p}}=[\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{n}]$, a token $\mathbf{p}_{i}$ for each of the parts specified in the part text prompt. We denote the $\hat{\mathbf{b}}$ to denote the base text prompt and $\hat{\mathbf{p}}$ to denote the part prompt. For the time $t\leq T_{th}$ (by default, we use $T_{th}\approx T/2$), we denoise the U-Net in the object region by denoising it with both the combination of the Part Prompt output and Base Prompt output. Due to this, the Part-Based U-Net gets the information regarding various parts in the object (Fig. 3). We now mathematically define the output noise $\epsilon_{t}$ for the diffusion process with Part-Diffusion below:
$\epsilon_{t}=\alpha\mathcal{M}_{o}\odot D(x_{t},\hat{\mathbf{p}},t)+(1-\alpha\mathcal{M}_{o})\odot D(x_{t},\hat{\mathbf{b}},t)$ (1)
@@
here, the $\alpha$ is the hyper-parameter controlling the strength of the part prompt diffusion output, and $D$ is the output of the pre-trained U-Net of a text-to-image diffusion model. The above denoising process is followed for t steps until the last step to produce a base image corresponding to the given prompt $\hat{b}$. Keeping a high $T_{th}$ and low $\alpha$, makes minimal changes in output as if the denoising diffusion was done with original prompt $\hat{b}$ (See Suppl. Sec.). With this part denoising of object, we obtain attention masks, from which we extract the localization information of part $\mathbf{p}_{t}$. We now describe the process of obtaining the part masks.
[…]
@@
Here $\delta$ is a hyperparameter, and $K$ is the number of parts. is This condition is robust in finding the localized parts (See Sec. 5 for ablation). For the parts $\mathbf{p}_{i}$, which are localized, we normalize the cross-attention map:
$\mathbf{\hat{m}}_{j}=L(j)\frac{\mathbf{\hat{m}}_{j}-\min(\mathbf{\hat{m}}_{j})}{\max(\mathbf{\hat{m}}_{j})-\min(\mathbf{\hat{m}}_{j})}+(1-L(j))\;\hat{\mathbf{m}}_{j}.$
We follow a dot-product-based protocol to assign each K cluster in the self-attention masks to a part, unlike the average attention protocol in the previous works [37, 18]. We find that dot product of normalized cross attention scores $\hat{\mathbf{m}}_{j}$ for each token with self-attention masks $\mathbf{\hat{M}}_{j}$ works better, as the attention maps in Part U-Net are noisy. Still, they are often correct for the regions that are localized in one specific area only (Suppl. Fig. 14). Hence, the dot product protocol favors those maps that are only localized in some areas of the image and don’t have high attention values across all parts of the object. After obtaining the dot product scores for each part, we assign $\hat{\mathbf{M}}_{j}$ to $\mathbf{p}_{i}$ with the highest scores. The mask for the part $\mathbf{p}_{i}$ is finally given as the following:
@@
Here $\epsilon$ is the hyperparameter which controls the minimum similarity required between the attention mask and the part token. We combine the additional attention masks unassigned to any token, and name them as the background (other) token $\mathbf{M}_{b}$.